OpenStudy (lindsay):

Consider a 10-question test in which each question can be answered either correctly or incorrectly. How many different ways are there to answer the question so that at least eight are correct? The answer is 56..how??

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (lindsay):

thanks

OpenStudy (lindsay):

I know you are supposed to use combinations to solve this, but I am stuck, please help

OpenStudy (anonymous):

as the question says atleast 8 it implies either 8 or 9 or 10 questions are correct....... => 10c8 ( for 8 correct) + 10c9(for 9 ) + 10c10 (for 10) = 45 + 10+1=56